I need help finding the slope and the y intercept of 4y-2x=12
Answer by mananth(16946)

4y-2x=12
4y=2x+12
/4
y=(1/2)x+3
compare wit the equation y = mx+b
m=slope = 1/2
y intercept = 3
y intercept point = (0,3)